Circular No. 6765

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF URANUS Orbital computations by B. G. Marsden and G. V. Williams, Harvard-

Smithsonian Center for Astrophysics, suggest that the observations of S/1997 U 2 (cf. IAUC 6764) are incompatible with a heliocentric orbit and that the best-fitting uranicentric orbits are quite eccentric, with a retrograde solution being significantly preferable to a direct solution. The poorer distribution of observations of S/1997 U 1 renders this situation very much less conclusive. Uranicentric orbital elements (referred to the ecliptic) and ephemerides follow for S/1997 U 1 and S/1997 U 2, respectively, the orbit for the former being arbitrarily adopted as retrograde and the same size (a = 0.038356 AU, n^{o} = 0.86762, P = 415 days) as that for the latter. Absolute magnitudes are H ∼ 9.0 and 7.5, respectively; for an albedo of 0.07, the radii of the objects would be ∼ 40 and 80 km.

T = 1997 Feb. 7.812 TT  ω = 13.^{o}242
e = 0.16014  Ω = 182.292 }2000.0
q = 0.03221 AU  i = 137.845
1997TT α_{2000} δ_{2000} Δ r_{U} ϵ β V
Oct. 9 20^{h}29.^{m}43 −19^{o}43.^{′}0 19.521 0.044 109.^{o}1 2.^{o}7 22.3
19 20 29.37 −19 43.4 19.689 0.043  99.1 2.8 22.3
29 20 29.65 −19 42.5 19.860 0.043  89.3 2.8 22.3
Nov. 8 20 30.28 −19 40.3 20.031 0.042  79.4 2.8 22.3
18 20 31.24 −19 36.8 20.196 0.042  69.6 2.6 22.3
28 20 32.51 −19 32.2 20.350 0.041  59.8 2.4 22.3
Dec. 8 20 34.06 −19 26.4 20.488 0.040  50.0 2.2 22.3
T = 1998 Apr. 2.675 TT  ω = 232.^{o}919
e = 0.52112  Ω = 210.613 }2000.0
q = 0.01837 AU  i = 149.423
1997TT α_{2000} δ_{2000} Δ r_{U} ϵ β V
Oct. 9 20^{h}28.^{m}79 −19^{o}37.^{′}6 19.440 0.057 108.^{o}9 2.^{o}8 20.7
19 20 28.82 −19 37.2 19.607 0.057  99.0 2.9 20.7
29 20 29.20 −19 35.6 19.778 0.056  89.2 2.9 20.7
Nov. 8 20 29.93 −19 32.8 19.950 0.055  79.3 2.8 20.8
18 20 30.99 −19 28.8 20.116 0.053  69.5 2.7 20.8
28 20 32.36 −19 23.8 20.273 0.052  59.8 2.5 20.8
Dec. 8 20 34.02 −19 17.8 20.416 0.050  50.0 2.2 20.8
1997 October 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden